OpenStudy (anonymous):

4(x - 6) = -2(x - 3)

OpenStudy (confluxepic):

What are you trying to find?

OpenStudy (anonymous):

simplify each side so that 4x-24=-2x+6

OpenStudy (anonymous):

combine like terms and divide, and you have x

OpenStudy (anonymous):

does x = 15? That's what I got

OpenStudy (anonymous):

Give me a moment to work it out.

OpenStudy (anonymous):

It should

OpenStudy (anonymous):

thank you so much for your help I hope it's right!

OpenStudy (anonymous):

If its not 15 is 5.

OpenStudy (anonymous):

* its

OpenStudy (anonymous):

I think it's 5 actually thanks for catching that! x = 15 4(15 - 6) = -2(15 - 3) 4 * 9 = -2 * 12 36 ≠ -24 but if x = 5 4(5 - 6) = -2(5 - 3) 4 * (-1) = -2 * 2 -4 = -4

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

I had to do the same thing you just did lol.

OpenStudy (anonymous):

thanks for catching that :)

OpenStudy (anonymous):

No prob :P